# Saner: Annotation-free Societal Attribute Neutralizer for Debiasing CLIP

### v1

@@ Abstract
Large-scale vision-language models, such as CLIP, are known to contain harmful societal bias regarding protected attributes (e.g., gender and age). In this paper, we aim to address the problems of societal bias in CLIP. Although previous studies have proposed to debias societal bias through adversarial learning or test-time projecting, our comprehensive study of these works identifies two critical limitations: 1) loss of attribute information when it is explicitly disclosed in the input and 2) use of the attribute annotations during debiasing process. To mitigate societal bias in CLIP and overcome these limitations simultaneously, we introduce a simple-yet-effective debiasing method called SANER (societal attribute neutralizer) that eliminates attribute information from CLIP text features only of attribute-neutral descriptions. Experimental results show that SANER, which does not require attribute annotations and preserves original information for attribute-specific descriptions, demonstrates superior debiasing ability than the existing methods.
Keywords:
Societal bias CLIP Debiasing Fairness
@@ 1 Introduction
Large-scale vision-language models (VLMs), such as CLIP [37], have demonstrated a remarkable capability in multi-modal understanding [32, 50] and generation [39, 48, 57], being trained with million-scale image-text pairs. Utilizing these VLMs, recent vision models have achieved significant performance enhancements across a wide range of computer vision tasks (e.g., captioning [15, 33, 58, 28, 56] and object detection [27, 60]), without the necessity for task-specific training [45].
Despite the success, several works have identified societal bias regarding demographic attributes, such as gender and age, in these VLMs [54, 18, 36, 5, 21, 1, 55], potentially causing unfair or prejudicial decisions by models. Hall et al. [20] conducted audits on performance disparity, particularly with respect to gender, and revealed gender-dependency of the CLIP performance. Qiu et al. [36] also demonstrated that adopting CLIP for caption evaluation tends to favor gender-stereotypical sentences (e.g., preferring “A woman is cooking” over “A man is cooking” for images depicting men), highlighting the inherent gender bias. These findings underscore the importance of addressing bias in VLMs.
Some studies have proposed to mitigate societal bias in VLMs [4, 43, 11, 9]. Adversarial debiasing [4, 43, 11] fine-tunes CLIP to lessen leakage of protected attributes [^1] We refer to any demographic variables, like age and gender, as protected attribute (or attribute in short), based on which a model’s decisions should not be made. into the features, while projection-based debiasing [9] removes the protected attribute encoded in CLIP features in the inference phase. Our holistic review of these pioneering works (Sec. 3), though, identifies the following potential disadvantages or controversies in their design choices.
Loss of attribute information explicitly disclosed in the input. Some methods aim to completely remove attribute information by decorrelating the attribute and the features [11] or by squashing the subspace associated with the attribute [9], even when the attribute is explicitly disclosed. This choice can limit the generalizability of a VLM’s features to a spectrum of downstream tasks (e.g., Stable Diffusion [39] generates male images for text prompt “a female doctor” when encoded with debiased CLIP as shown in Fig. 1 (a) and Sec. 6.2), while it works for attribute-agnostic downstream tasks [26, 35].
Use of the attribute annotations. Adversarial debiasing methods [4, 43, 11] require protected attribute annotations, as provided in FairFace [24], for fine-tuning (Fig. 1 (b)). Datasets with attribute annotations are still scarce, perhaps partly because the annotation process needs ethical considerations [2], limiting their applicability and generalizability. This dataset scarcity also causes the limited diversity of images and text descriptions used for fine-tuning the VLM, potentially inducing overfitting.
@@
Contribution. Thanks to our annotation-free debiasing pipeline, SANER is designed to be compatible with any dataset of image-text pairs, such as COCO [29]. This provides denser guidance for training the debiasing layer compared to the existing methods. Experiments on both discriminative and generative tasks (i.e., text-to-image retrieval [19] and text-to-image generation [39]) show that SANER can mitigate gender, age, and racial biases of CLIP. Moreover, we demonstrate that SANER outperforms the existing methods [4, 9], showing that SANER leads to less attribute-dependency of the downstream performance while overcoming the limitations in existing methods.
2 Related Work
Societal bias in VLMs. With the growing application of VLMs, such as CLIP [37], across various tasks, there has been increasing awareness of their inherent social biases [20, 12, 51, 40, 55, 41, 46]. Hall et al. [20] examined gender bias in CLIP, uncovering significant discrepancies in object recognition performance based on the gender depicted in images. For example, the accuracy in predicting an umbrella when it appears with women is much better than with men.
The existence of such societal bias poses a substantial risk of perpetuating and exacerbating discrimination against historically marginalized groups [5, 36, 47, 53, 22]. Birhane et al. [5] demonstrated that adopting CLIP-based filtering in the dataset creation causes a risk of selecting stereotypical images regarding specific demographic subgroups. For instance, for an image of a female astronaut, the caption “This is a photograph of a smiling housewife” is favored over “This is a portrait of an astronaut” by CLIP, resulting in datasets that contain a number of image-caption pairs with harmful stereotypical descriptions.
Debiasing societal bias in CLIP. Recognizing the problem of bias in CLIP, some recent research has concentrated on mitigating societal bias in CLIP [4, 43, 11, 9]. These debiasing efforts intend to remove attribute information from CLIP features, resulting in less biased prediction outcomes. For example, Berg et al. [4] proposed a prompt tuning-based debiasing method based on an adversarial learning paradigm, minimizing the encoding of protected attribute information by the CLIP text encoder. However, in Sec. 3, we identify that these methods contain critical limitations that pose challenges in utilizing various datasets for debiasing and applying debiased models to a wide range of downstream tasks.
@@ 3 Review: Existing Debiasing Methods
Several debiasing approaches for CLIP have been introduced, broadly categorized into two main types: adversarial debiasing [4, 43, 11] and projection-based debiasing [9]. This section conducts an in-depth analysis of these existing debiasing strategies, highlighting their respective limitations.
@@ 3.1 Adversarial debiasing
Adversarial debiasing [4, 43, 11] aims to eliminate protected attribute information in the CLIP features. Specifically, an adversarial classifier is employed to predict and remove protected attribute $a$ from CLIP features.
@@
Let $s(v)\in\mathbb{R}^{|\mathcal{C}|}$ denote a vector, each of whose elements is the similarity score $s_{c}(v)$ for a concept in $\mathcal{C}$. Due to the attribute non-specificity of concepts in $\mathcal{C}$, $s(v)$ should not correlate with attribute $a$ of $v$. However, the CLIP text encoder can embed $a$ into $s(v)$ due to bias, allowing an attribute classifier to predict $a$. We denote the probability of being $a$ given $s(v)$ (or a prediction score of the attribute classifier) by $m_{a}(s(v))$. Prompt tuning-based debiasing [4] uses $m_{a}$ for adversarial loss, given by
@@
The experiments [4] showed that this method could effectively reduce attribute leakage through $f_{\text{t}}(t)$, but the limited number of concepts [^3] Their experiments used $10$ concepts. may limit downstream tasks that enjoy the debiased features because as $l$ is learned only through a sparse set of concepts. Additionally, attribute annotations are necessary for the adversarial loss, resulting in the exclusive use of face-centric image datasets (e.g., FairFace [24]) as $\mathcal{D}$.
Additive Residual Learner (ARL) [43] is designed to remove attribute information from CLIP image features. This method assumes that a debiasing layer [^4] A fully-connected layer is used. $r$ can identify a vector to additively amend attribute-neutral image feature vector $\delta(v)$, i.e.,
@@
Similarly to [4], an adversarial classifier is trained to predict $a$ from $\delta$ with adversarial loss
@@
The reconstruction loss between $f_{\text{v}}(v)$ and $\delta(v)$ regularizes the training to maintain the original features.
This method shares a common limitation with prompt tuning-based debiasing [4], notably requiring attribute annotations. Another limitation is that it tries to remove attribute features even when attributes of people in images are explicitly disclosed (i.e., when the person is depicted in an image). Consequently, debiased CLIP is ignorant of protected attributes.
Mapper [11] aims to reduce spurious correlations between protected attributes $a$ and task label $d$ in $\mathcal{D}$. It applies mappings $f_{\text{v}}^{\prime}$ and $f_{\text{t}}^{\prime}$ to image and text features, respectively, as $x_{\text{v}}(v)=f_{\text{v}}^{\prime}(f_{\text{v}}(v))$ and $x_{\text{t}}(t)=f_{\text{t}}^{\prime}(f_{\text{t}}(t))$ for mitigating dependence on $a$. The adversarial loss is computed using a dependence measure $\text{Dep}(\cdot,\cdot)$ to quantify statistical dependence between features as:
@@
Similar to [4, 43], Mapper relies on attribute annotations. Moreover, it is designed only to address the spurious correlations between the attribute and task labels for a specific task but not for different tasks.
@@ 3.2 Projection-based debiasing
where $I$ is the identity matrix. $P$ can project a CLIP text feature vector $f_{\text{t}}(t)$ for a text description $t$ into the orthogonal complement by $Pf_{\text{t}}(t)$.
@@ 3.3 Summary of the challenges
The existing debiasing methods, including adversarial and projection-based, reveal several challenges: 1) Loss of attribute information (ARL and projection-based) even with explicit attribute description narrows down the utility of the debiased CLIP. For example, a text-to-image generative model with gender-debiased CLIP features may not properly depict explicitly specified gender (as
shown in Sec. 6.2), 2) Dependency on attribute annotations (prompt tuning, ARL, and Mapper) constrains the range of datasets that can be utilized, often necessitating the use of face-centric image datasets (e.g., FairFace), as opposed to more diverse, natural image datasets (e.g., COCO [29]).
@@ 4 Societal Attribute Neutralizer (SANER)
Figure 2: An overview of SANER, exemplified by binary gender.
@@ 4.1 Attribute neutralization
We first modify text description $t\in\mathcal{D}$ that contain person-related words, [^5] The person-related words include any words relevant to a person (e.g., person, girl, man). The complete list is available in the supplementary material. to remove attribute-specific words. Taking binary gender as a protected attribute, [^6] Following prior research [4, 43, 11, 9, 59, 52, 6], we focus on the binary gender but recognize the importance of inclusivity. SANER applies to non-binary genders. i.e., $\mathcal{A}=\{\texttt{female},\texttt{male}\}$, as example, the text description
@@
contains attribute information (i.e., woman). We replace attribute-specific terms [^7] We use gender words defined in [23]. with the attribute-neutral ones to obtain an attribute-neutral text:
@@
where $\xi_{\text{n}}$ denotes a function for attribute neutralization. Neutralization can be done for other attributes, such as age. [^8] Examples for the race attribute are in the supplementary material. We remove age-specific terms [^9] We define age-specific terms. The list is in the supplementary material. (e.g., young and senior) in text descriptions, for instance, “A young woman is eating salad” $\rightarrow$ “A woman is eating salad”. In contrast to the previous approach [9], which is optimized not to predict the attribute information from the original description $t$, we target the attribute-neutral descriptions $\xi_{\text{n}}(t)$ to preserve the attribute information in the features of attribute-specific descriptions.
@@ 4.2 Feature modification
CLIP text features $z(t)=f_{\text{t}}(\xi_{\text{n}}(t))$ after attribute neutralization can still convey the protected attribute information due to CLIP’s bias. To remove such bias, we append a learnable debiasing layer $r$ on top of $f_{\text{t}}$, inspired by recent CLIP fine-tuning techniques [43, 17]. Neutralized $t$’s debiased feature $h(t)$ is given by
$h(t)=z(t)+r(z(t)).$ (8)
@@ 4.3 Attribute annotation-free debiasing loss
To train $r$ to extract attribute information from CLIP features without attribute annotations, we generate a set $\mathcal{T}$ of attribute-specific descriptions for $t\in\mathcal{D}$ and for $g\in\mathcal{A}$, i.e., $\mathcal{T}=\{\xi_{g}(t)|t\in\mathcal{D},g\in\mathcal{A}\}$, where $\xi_{g}(t)$ generates a description specific to attribute group $g$ from $t$. For the binary gender case, this involves generating descriptions with female- and male-specific words. For instance, from the text description, “A woman is eating salad.”, we generate two sentences with female and male attributes:
@@
The debiasing loss trains $r$ such that $h(t)$ is equidistant from $f_{\text{t}}(\xi_{g}(t))$ for all attribute groups in $\mathcal{A}$, ensuring an impartial representation across the spectrum of attribute groups. We implement this loss as the standard deviation of the cosine similarity between $h(t)$ and $f_{\text{t}}(\xi_{g}(t))$. Let $s_{g}(t)$ denote the similarity, i.e.,
$s_{g}(t)=\frac{h(t)^{\top}f_{\text{t}}(\xi_{g}(t))}{\|h(t)\|\;\|f_{\text{t}}(\xi_{g}(t))\|}.$ (9)
@@
where $\bar{s}(t)=\sum_{g\in\mathcal{A}}s_{g}(t)/|\mathcal{A}|$. A lower standard deviation means $s_{g}$ is close to $\bar{s}$, leading to $h(t)$ being equidistant to $f_{\text{t}}(\xi_{g}(t))$ for all $g\in\mathcal{A}$. Notably, this debiasing loss can be computed without attribute annotations.

### v3

@@ Abstract
Large-scale vision-language models, such as CLIP, are known to contain societal bias regarding protected attributes (e.g., gender, age). This paper aims to address the problems of societal bias in CLIP. Although previous studies have proposed to debias societal bias through adversarial learning or test-time projecting, our comprehensive study of these works identifies two critical limitations: 1) loss of attribute information when it is explicitly disclosed in the input and 2) use of the attribute annotations during debiasing process. To mitigate societal bias in CLIP and overcome these limitations simultaneously, we introduce a simple-yet-effective debiasing method called SANER (societal attribute neutralizer) that eliminates attribute information from CLIP text features only of attribute-neutral descriptions. Experimental results show that SANER, which does not require attribute annotations and preserves original information for attribute-specific descriptions, demonstrates superior debiasing ability than the existing methods. Additionally, we observe that SANER does not require retraining CLIP from scratch with the original dataset. Moreover, the debiased model can be directly applied to the text-to-image generation model by simply replacing the text encoder. [^1] Project page: https://rebnej.github.io/saner-clip.github.io/
@@ 1 Introduction
Large-scale vision-language models (VLMs), such as CLIP (Radford et al., 2021), have demonstrated a remarkable capability in multi-modal understanding (Lüddecke & Ecker, 2022; Tewel et al., 2022) and generation (Rombach et al., 2022; Tao et al., 2023; Yamazaki et al., 2023), being trained with million-scale image-text pairs. Utilizing these VLMs, recent vision models have achieved significant performance enhancements across a wide range of computer vision tasks (e.g., captioning (Mokady et al., 2021; Li et al., 2023b; Yamazaki et al., 2022) and object detection (Li et al., 2022; Zhong et al., 2022)), without the necessity for task-specific training (Shen et al., 2022).
Despite the success, several works have identified societal bias regarding demographic attributes, such as gender and age, in these VLMs (Wolfe & Caliskan, 2022; Hausladen et al., 2024; Alabdulmohsin et al., 2024), potentially causing unfair or prejudicial decisions by models. Hall et al. (2023) conducted audits on performance disparity, particularly with respect to gender, and revealed gender-dependency of the CLIP performance. Qiu et al. (2023) also demonstrated that adopting CLIP for caption evaluation tends to favor gender-stereotypical sentences (e.g., preferring “A woman is cooking” over “A man is cooking” for images depicting men), highlighting the inherent gender bias. These findings underscore the importance of addressing bias in VLMs.
Some studies have proposed mitigating societal bias in VLMs (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024; Chuang et al., 2023). Adversarial debiasing (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024) fine-tunes CLIP to lessen leakage of protected attributes [^2] We refer to any demographic variables, like age and gender, as protected attribute (or attribute in short), based on which a model’s decisions should not be made. into the features, while projection-based debiasing (Chuang et al., 2023) removes the protected attribute encoded in CLIP features at the inference phase. Our holistic review of these pioneering works (Sec. 2), though, identifies the following potential drawbacks or controversies in their design choices.
Loss of attribute information explicitly disclosed in the input. Some methods aim to completely remove attribute information by decorrelating the attribute and the features (Dehdashtian et al., 2024) or by squashing the subspace associated with the attribute (Chuang et al., 2023), even when the attribute is explicitly disclosed. This choice can limit the generalizability of a VLM’s features to a spectrum of downstream tasks (e.g., Stable Diffusion (Rombach et al., 2022) generates male images for text prompt “a female doctor” when encoded with debiased CLIP as shown in Fig. 1 (a) and Sec. 5.2), while it works for attribute-agnostic downstream tasks (Krause et al., 2013).
Use of the attribute annotations. Adversarial debiasing methods (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024) require protected attribute annotations, as provided in FairFace (Karkkainen & Joo, 2021), for fine-tuning (Fig. 1 (b)). Datasets with attribute annotations are still scarce, partly because the annotation process needs ethical considerations (Andrews et al., 2023), limiting their applicability. This dataset scarcity also causes the limited diversity of images and text descriptions used for fine-tuning the VLM, potentially inducing overfitting.
@@
Contribution. Thanks to our annotation-free debiasing pipeline, SANER is designed to be compatible with any dataset of image-text pairs, such as COCO (Lin et al., 2014). This provides denser guidance for training the debiasing layer compared to the existing methods. Moreover, SANER does not require retraining the CLIP model itself, accessing its original training data, or retraining downstream tasks (e.g., text-to-image generation) when applying the debiased CLIP. Experiments on both discriminative and generative tasks (i.e., text-to-image retrieval (Geyik et al., 2019) and text-to-image generation (Rombach et al., 2022)) show that SANER can mitigate gender, age, and racial biases of CLIP. Moreover, we demonstrate that SANER outperforms the existing methods (Berg et al., 2022; Chuang et al., 2023), showing that SANER leads to less attribute-dependency of the downstream performance while overcoming the limitations in existing methods.
@@ 2 Review: Existing Debiasing Methods
Several debiasing approaches for CLIP have been introduced, broadly categorized into two main types: adversarial debiasing (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024) and projection-based debiasing (Chuang et al., 2023). This section conducts an in-depth analysis of these existing debiasing strategies, highlighting their respective limitations.
@@ 2.1 Adversarial debiasing
Adversarial debiasing (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024) aims to eliminate protected attribute information in the CLIP features. Specifically, an adversarial classifier is employed to predict and remove protected attribute $a$ from CLIP features.
@@
Let $s(v)\in\mathbb{R}^{|\mathcal{C}|}$ denote a vector, each of whose elements is the similarity score $s_{c}(v)$ for a concept in $\mathcal{C}$. Due to the attribute non-specificity of concepts in $\mathcal{C}$, $s(v)$ should not correlate with attribute $a$ of $v$. However, the CLIP text encoder can embed $a$ into $s(v)$ due to bias, allowing an attribute classifier to predict $a$. We denote the probability of being $a$ given $s(v)$ (or a prediction score of the attribute classifier) by $m_{a}(s(v))$. Prompt tuning-based debiasing (Berg et al., 2022) uses $m_{a}$ for adversarial loss, given by
@@
The experiments (Berg et al., 2022) showed that this method could effectively reduce attribute leakage through $f_{\text{t}}(t)$, but the limited number of concepts [^4] Their experiments used $10$ concepts. may limit downstream tasks that enjoy the debiased features because as $l$ is learned only through a sparse set of concepts. Additionally, attribute annotations are necessary for the adversarial loss, resulting in the exclusive use of face-centric image datasets (e.g., FairFace (Karkkainen & Joo, 2021)) as $\mathcal{D}$.
Additive Residual Learner (ARL) (Seth et al., 2023) is designed to remove attribute information from CLIP image features. This method assumes that a debiasing layer [^5] A fully-connected layer is used. $r$ can identify a vector to additively amend attribute-neutral image feature vector $\delta(v)$, i.e.,
@@
Similarly to (Berg et al., 2022), an adversarial classifier is trained to predict $a$ from $\delta$ with adversarial loss
@@
The reconstruction loss between $f_{\text{v}}(v)$ and $\delta(v)$ regularizes training to preserve the original features.
This method shares a common limitation with prompt tuning-based debiasing (Berg et al., 2022), notably requiring attribute annotations. Another limitation is that it tries to remove attribute features even when attributes of people in images are explicitly disclosed (i.e., when the person is depicted in an image). Consequently, debiased CLIP is ignorant of protected attributes.
Mapper (Dehdashtian et al., 2024) aims to reduce spurious correlations between attributes $a$ and task label $d$ in $\mathcal{D}$. It applies mappings $f_{\text{v}}^{\prime}$ and $f_{\text{t}}^{\prime}$ to image and text features, respectively, as $x_{\text{v}}(v)=f_{\text{v}}^{\prime}(f_{\text{v}}(v))$ and $x_{\text{t}}(t)=f_{\text{t}}^{\prime}(f_{\text{t}}(t))$ for mitigating dependence on $a$. The adversarial loss is computed using a dependence measure $\text{Dep}(\cdot,\cdot)$ to quantify statistical dependence between features as:
@@
Similar to (Berg et al., 2022; Seth et al., 2023), Mapper relies on attribute annotations. Moreover, it is designed only to address the spurious correlations between the attribute and task labels for a specific task but not for different tasks.
@@ 2.2 Projection-based debiasing
where $I$ is the identity matrix. $P$ can project a CLIP text feature vector $f_{\text{t}}(t)$ for a text description $t$ into the orthogonal complement by $Pf_{\text{t}}(t)$. This process removes attribute information by projecting features into the space orthogonal to attribute-specific directions.
@@ 2.3 Summary of the challenges
The existing debiasing methods, including adversarial and projection-based, reveal several challenges: 1) Loss of attribute information (ARL and projection-based) even with explicit attribute description narrows down the utility of the debiased CLIP. For example, a text-to-image generative model with gender-debiased CLIP features may not properly depict explicitly specified gender (as shown in Sec. 5.2), 2) Dependency on attribute annotations (prompt tuning, ARL, and Mapper) constrains the range of datasets that can be utilized, often necessitating the use of face-centric image datasets (e.g., FairFace), as opposed to more diverse, natural image datasets (e.g., COCO).
@@ 3.1 Attribute neutralization
We first modify text description $t\in\mathcal{D}$ that contain person-related words, [^6] Person-related words encompass terms that reference individuals (e.g., person, girl, man). The complete list is in the appendix. to remove attribute-specific words. Taking binary gender as a protected attribute, [^7] Following prior research (Berg et al., 2022; Seth et al., 2023; Dehdashtian et al., 2024; Chuang et al., 2023; Zhao et al., 2017; Garcia et al., 2023; Burns et al., 2018), we focus on the binary gender but recognize the importance of inclusivity. SANER applies to non-binary genders. i.e., $\mathcal{A}=\{\texttt{female},\texttt{male}\}$, as example, the text description
@@
contains attribute information (i.e., woman). We replace attribute-specific terms [^8] We use gender words defined in (Hirota et al., 2023). with the attribute-neutral ones to obtain an attribute-neutral text:
@@
where $\xi_{\text{n}}$ denotes a function for attribute neutralization. Neutralization can be done for other attributes, such as age. [^9] Examples for the race attribute are in the appendix. We remove age-specific terms [^10] We define age-specific terms. The list is in the appendix. (e.g., young and senior) in text descriptions, for instance, “A young woman is eating salad” $\rightarrow$ “A woman is eating salad”. In contrast to the previous approach (Chuang et al., 2023), which is optimized not to predict the attribute information from the original description $t$, we target the attribute-neutral descriptions $\xi_{\text{n}}(t)$ to preserve the attribute information in the features of attribute-specific descriptions.
Figure 2: An overview of SANER, exemplified by binary gender. SANER neutralizes attribute-specific text (e.g., “woman” $\rightarrow$ “person”), modifies features via debiasing layer, and uses three losses for debiasing: $\mathcal{L}_{\text{deb}}$ for attribute neutralization, $\mathcal{L}_{\text{recon}}$ for feature preservation, and $\mathcal{L}_{\text{cont}}$ for image-text alignment.
@@ 3.2 Feature modification
CLIP text features $z(\xi_{\text{n}}(t))=f_{\text{t}}(\xi_{\text{n}}(t))$ after attribute neutralization can still convey the protected attribute information due to CLIP’s bias. To remove such bias, we append a learnable debiasing layer $r$ on top of $f_{\text{t}}$, inspired by recent CLIP fine-tuning techniques (Seth et al., 2023; Gao et al., 2024). Neutralized $t$’s debiased feature $h(\xi_{\text{n}}(t))$ is given by
$h(\xi_{\text{n}}(t))=z(\xi_{\text{n}}(t))+r(z(\xi_{\text{n}}(t))).$ (8)
@@ 3.3 Attribute annotation-free debiasing loss
To train $r$ to extract attribute information from CLIP features without attribute annotations, we create a set $\mathcal{T}$ of attribute-specific descriptions for $t\in\mathcal{D}$ and for $g\in\mathcal{A}$, i.e., $\mathcal{T}=\{\xi_{g}(t)|t\in\mathcal{D},g\in\mathcal{A}\}$, where $\xi_{g}(t)$ generates a description for each attribute group $g$ from $t$. For binary gender, this involves generating descriptions with female- and male-specific words. For instance, from the text description, “A woman is eating salad.”, we generate two sentences with female and male attributes:
@@
The debiasing loss trains $r$ such that $h(\xi_{\text{n}}(t))$ is equidistant from $f_{\text{t}}(\xi_{g}(t))$ for all attribute groups in $\mathcal{A}$, ensuring an impartial representation across the spectrum of attribute groups. We implement this loss as the standard deviation of the cosine similarity between $h(\xi_{\text{n}}(t))$ and $f_{\text{t}}(\xi_{g}(t))$. Let $s_{g}(t)$ denote the similarity, i.e.,
$s_{g}(t)=\frac{h(\xi_{\text{n}}(t))^{\top}f_{\text{t}}(\xi_{g}(t))}{\|h(\xi_{\text{n}}(t))\|\;\|f_{\text{t}}(\xi_{g}(t))\|}.$ (9)
@@
where $\bar{s}(t)=\sum_{g\in\mathcal{A}}s_{g}(t)/|\mathcal{A}|$. A lower standard deviation means $s_{g}$ is close to $\bar{s}$, leading to $h(\xi_{\text{n}}(t))$ being equidistant to $f_{\text{t}}(\xi_{g}(t))$ for all $g\in\mathcal{A}$. Notably, this debiasing loss can be computed without attribute annotations.